# DISC: Plug-and-Play Decoding Intervention with Similarity of Characters for Chinese Spelling Check

### v1

@@ 1 Introduction
Given an input sentence, Chinese spelling check (CSC) aims to detect incorrect characters and modify each into an correct character Yu and Li (2014); Xu et al. (2021). Table 1 gives two examples. Spelling errors degrade reading efficiency, and sometimes even lead to misunderstanding. The authority or attitude of the writer may be doubted if their document contains simple spelling errors. Moreover, spelling errors substantially hurt the performance of subsequent NLP models.
@@
Previous works employ confusion sets to leverage such similarities among characters Yeh et al. (2013); Huang et al. (2014); Xie et al. (2015); Cheng et al. (2020); Huang et al. (2023). Formally, a confusion set is denoted as $\mathcal{C}=\{(c^{1}_{i},c^{2}_{i})\}_{i=1}^{M}$, where $(c^{1}_{i},c^{2}_{i})$ represents a pair of characters and means that $c^{1}_{i}$ may be mistakenly replaced by $c^{2}_{i}$ in real texts.
@@
As a replacement for confusion sets, we propose a lightweight plug-and-play DISC (decoding intervention with similarity of characters) module. DISC derives probability-based similarities among characters in both phonetics and glyph, and uses the probabilities to intervene in the decoding process. Similar to the confusion set, our DISC aims to enhance the model’s precision. However, for datasets that lack or have no in-domain training data, DISC may result in under-corrections due to the model’s conservative predictions, leading to unstable recall. To address this, we propose a copy-punishment solution to balance precision and recall.
@@
Experiments and analyses on popular benchmark datasets, i.e., SIGHANs, ECSpell, and LEMON, demonstrate that our DISC module can significantly enhance the error correction performance of CSC models. This improvement does not require additional training costs and only slightly affects the decoding efficiency of the model. We release our code at https://anonymous.4open.science/r/simple-DISC.
@@ 4.1 Datasets
However, many previous studies have pointed out that the SIGHAN datasets may not represent real-world CSC tasks, as they are derived from Chinese learner texts. To address this limitation, we also conduct experiments on the ECSpell Lv et al. (2023) and LEMON Wu et al. (2023) datasets, which are derived from Chinese native-speaker (CNS) texts and encompass a wide range of domains. It is worth noting that LEMON does not have a dedicated training set, making it an excellent test set for evaluating a model’s generalization ability.
@@ 4.2 Baseline Models
In the era of LLMs, researchers have begun using LLMs to explore the CSC field. We present the results of representative LLMs on certain benchmarks for comparison, including the top-performing GPT series in terms of overall capability: GPT3.5 and GPT4, as well as some results on open-source LLMs in the Chinese NLP community from previous work, such as finetuned Baichuan2 Yang et al. (2023a).
@@
 + DISC 69.9 56.3 62.4 8.9
Table 3: Sentence-level performance of LLMs, ReLM, and ReLM + DISC on the test sets of ECSpell and LEMON. Results marked with $\star$” are from Liu et al. (2024).
@@ 5 Main Results
Table 4 illustrates the main results across SIGHAN benchmarks, demonstrating that the addition of the DISC module in the decoding process leads to notable improvements across all the compared models, reaching state-of-the-art performance. Specifically, ReaLiSe + DISC has increases of 0.1/1.1/0.6, SCOPE + DISC achieves lifts of 1.2/1.4/0.8, ReLM + DISC sees enhancements of 2.1/3.3/1.2 in correction-level $\mathrm{F}_{1}$ (C-$\mathrm{F}$) score on the SIGHAN13/14/15 test sets, respectively.
@@
Table 4.2 depicts that the incorporation of the DISC module into ReLM leads to substantial improvements of 1.1/3.7/0.7 C-$\mathrm{F}$ score compared to unenhanced ReLM in the LAW, MED and ODW domains, respectively. Table 4.2 also presents the performance of DISC on LEMON. After integrating the DISC module, the results of ReLM + DISC achieve notable improvements across all domains, and the average C-$\mathrm{F}$ has an increase of 3.2. This demonstrates that our DISC module yields stable and significant improvements in cross-domain CSC testing.
@@ 6 Discussion
Robustness of similarity hyperparameters.
For $\beta$, it shows the effect of the proportion of phonetic similarity in the total similarity on the model’s correction performance. The $\mathrm{F}_{1}$ score curve shows a clear trend of rising first and then decreasing, which indicates that phonetic and glyph similarities are complementary, with phonetic similarity being relatively more important than glyph similarity.
The balance between precision and recall.
We observe that this decrease in recall primarily occurs in the LEMON test set. The key distinction between these datasets is that LEMON contains less seen edit pairs in training data. As shown in Table 4, we calculate the proportion of edit pairs from each test set that appear in the training set. [^4] For LEMON and ECSpell, we conduct statistics using the pairs in the confusion set which is used to generate 34 million monolingual sentences. Models are prone to copy the source character when the target edit pair is not available in the training data.
For this type of test set, we can use a simple copy punishment combined with DISC, which reduces the probability of copying the original character during inference, to mitigate the decrease in recall. Detailed experimental results can be found in the Appendix C.
@@
Effectiveness of DISC module.
We degrade the DISC module to a simple confusion set constraint decoding strategy. We investigate two confusion sets: one derived from our similarity computation strategy [^5] We treat a character pair as confused if their similarity score exceeds 0.5. and another pre-existing one provided by Wang et al. (2018). The results are shown in the second part of Table 5. From the results, we can see that both confusion sets fail to consistently improve performance, indicating the strategy’s sensitivity to confusion set quality. The confusion set from Wang et al. (2018) improves SIGHAN15 by covering over 99% of its erroneous pairs but degrades performance on other test sets, highlighting the domain-specific limitations of such confusion sets.
Effectiveness of components of the DISC module.
@@ 7 Related Work
These works fall into three categories. (1) At only the inference phase. Wang et al. (2019); Bao et al. (2020) use the confusion set as constraints upon the search space, i.e., allowing the model to only consider characters in the confusion set.
(2) For data synthesis. Liu et al. (2021) use a confusion set $\mathcal{C}$ to synthesize data for training CSC models. For a given correct sentence, they randomly select a character (e.g., $c_{i}$), and replace it with an incorrect character (e.g., $c^{\prime}$). Only characters in the confusion set, i.e., $(c_{i},c^{\prime})\in\mathcal{C}$, are considered.
@@
Besides the use of confusion sets, there exist some works that directly utilize phonetic and glyph information to enhance CSC models. Liu et al. (2021); Li et al. (2022) add an extra task of predicting the phonetic of each input character. Xu et al. (2021) use GRU to encode Pinyin, and use CNN to encode glyphs (font pictures) for each input character, as extra character representations.
@@
Gou and Chen (2021) extract features such as probability and rank of the original character and the top 1 candidate character, and use SVM to determine whether the modification should be retained. Yin et al. (2024) first retrieve similar segments from the training set. Then, they intervene in the decoding process based on the segment (n-gram) similarity between the retrieved segments and the input. Lv et al. (2023) employ a word dictionary in the target domain to assist the decoding process.
@@ Appendix B Details of datasets
Table 6: Statistics of the datasets, including the number of sentences, the average length of sentences, and the number of errors.
@@ Appendix C Copy Punishment
Appendix D Impact on decoding efficiency
[…]
ReaLiSe 24.5
+ DISC 27.5 1.143$\times$
SCOPE 138.6
+ DISC 143.4 1.035$\times$
ReLM 12.7
+ DISC 12.8 1.010$\times$
Appendix E Prompt Examples
@@
Figure 4: Prompt templates used in GPT3.5 and GPT4.
In this work, we use the prompt-based method to activate the CSC ability of the GPT3.5 and GPT4. The prompt used for the baselines are shown in Figure 4.
Appendix F More Results
@@ D-P D-R D-F D-P D-R D-F D-P D-R D-F
In addition to the correction-level performance, which is our primary focus, we also present the detection-level experimental results of the CSC models, as shown in Table F. SCOPE + DR-CSC performs well at the detection level, primarily because they incorporate an additional detection network.
Since SIGHANs contains a lot of noise, we also conduct experiments on their revised versions (referred to as SIGHANs (rev.)) released by Yang et al. (2023b), which have undergone manual verification and error correction to ensure higher data quality. As shown in Table F and Table F, our DISC module also achieves consistent performance improvements on SIGHANs (rev.).

### v2

@@ 1 Introduction
Given an input sentence, Chinese spelling check (CSC) aims to detect incorrect characters and modify each into a correct character Yu and Li (2014); Xu et al. (2021). Table 1 gives two examples. Spelling errors degrade reading efficiency, and sometimes even lead to misunderstanding. The authority or attitude of the writer may be doubted if their document contains simple spelling errors. Moreover, spelling errors substantially hurt the performance of subsequent NLP models.
@@
Previous works employ confusion sets to leverage such similarities among characters Yeh et al. (2013); Huang et al. (2014); Xie et al. (2015); Cheng et al. (2020); Huang et al. (2023). Formally, a confusion set is denoted as $\mathcal{C}=\{(c^{1}_{i},c^{2}_{i})\}_{i=1}^{M}$, where each pair $(c^{1}_{i},c^{2}_{i})$ represents a pair of characters and means that $c^{1}_{i}$ may be mistakenly replaced by $c^{2}_{i}$ in real texts.
@@
As a replacement for confusion sets, we propose a lightweight plug-and-play DISC (decoding intervention with similarity of characters) module. DISC derives probability-based similarities among characters in both phonetics and glyph, and uses them to intervene in the decoding process. Similar to the confusion set, our DISC aims to enhance the model’s precision. However, for datasets that lack or have no in-domain training data, DISC may result in under-corrections due to the model’s conservative predictions, leading to unstable recall. To address this, we propose a copy-punishment solution to balance precision and recall.
@@
Experiments and analyses on popular benchmark datasets, i.e., SIGHANs, ECSpell, and LEMON, demonstrate that our DISC module can significantly enhance the error correction performance of CSC models. This improvement does not require additional training costs and only slightly affects the decoding efficiency of the model. We release our code at https://github.com/zhqiao-nlp/DISC.
@@ 4.1 Datasets
However, many previous studies have pointed out that the SIGHAN datasets may not represent real-world CSC tasks, as they are derived from Chinese learner texts. To address this limitation, we also conduct experiments on the ECSpell Lv et al. (2023) and LEMON Wu et al. (2023) datasets, which are derived from Chinese native-speaker (CNS) texts and encompass a wide range of domains. It is worth noting that LEMON does not have a dedicated training set, making it an excellent test set for evaluating a model’s generalization ability. Due to space constraints, we selected results from four domains for display and provided the average performance across all seven domains.
@@ 4.2 Baseline Models
In the era of LLMs, researchers have begun using LLMs to explore the CSC field. We present the results of representative LLMs on certain benchmarks for comparison, including the top-performing GPT series in terms of overall capability: GPT3.5 and GPT4, as well as some results on open-source LLMs in the Chinese NLP community from previous work, such as finetuned Baichuan2 Yang et al. (2023a). Specifically, we demonstrate the results of Baichuan2 on the ECSpell and LEMON test sets using supervised fine-tuning (SFT) Liu et al. (2024) and prompt-free training-free approach Zhou et al. (2024), representing the current SOTA performance.
@@
ReLM 58.1 45.1 50.6 11.7
 + DISC 73.7 42.5 53.7 4.5
Table 3: Sentence-level performance of LLMs, ReLM, and ReLM + DISC on the test sets of ECSpell and LEMON. Results marked with $\star$” are from Liu et al. (2024), and $\dagger$” are from Zhou et al. (2024).
@@ 5 Main Results
Table 4 illustrates the main results across SIGHAN benchmarks, demonstrating that the addition of the DISC module in the decoding process leads to notable improvements across all the compared models, reaching state-of-the-art performance. Specifically, ReaLiSe + DISC has increases of 0.1/1.1/0.6, SCOPE + DISC achieves lifts of 1.2/1.4/0.8, and ReLM + DISC sees enhancements of 2.1/3.3/1.2 in correction-level $\mathrm{F}_{1}$ (C-$\mathrm{F}$) score on the SIGHAN13/14/15 test sets, respectively.
@@
Table 4.2 depicts that the incorporation of the DISC module into ReLM leads to substantial improvements of 1.1/3.7/0.7 C-$\mathrm{F}$ score compared to unenhanced ReLM in the LAW, MED and ODW domains, respectively.
Table 4.2 also presents the performance of DISC on LEMON. After integrating the DISC module, the results of ReLM + DISC achieve notable improvements across all domains, and the average C-$\mathrm{F}$ has an increase of 3.1. This demonstrates that our DISC module yields stable and significant improvements in cross-domain CSC testing.
@@ 6 Discussion
Robustness of similarity hyperparameters.
For $\beta$, it shows the effect of the proportion of phonetic similarity in the total similarity on correction performance. The $\mathrm{F}_{1}$ score curve shows a clear trend of rising first and then decreasing, which indicates that phonetic and glyph similarities are complementary, with phonetic similarity being relatively more important than glyph similarity.
The balance between precision and recall.
We observe that this decrease in recall primarily occurs in the LEMON test set. The key distinction between these datasets is that LEMON contains less seen edit pairs in training data. As shown in Table 4, we calculate the proportion of edit pairs from each test set that appear in the training set. [^4] For LEMON, we conduct statistics using the pairs in the confusion set which is used to generate 34 million monolingual sentences. In ECSpell, the proportion of seen edit pairs is very low, yet the model performs well. This is due to data leakage, as we explain in Appendix E. Models are prone to copy the source character when the target edit pair is not available in the training data.
For this type of test set, we can use a simple copy punishment combined with DISC, which reduces the probability of copying the original character during inference, to mitigate the decrease in recall. Detailed experimental results can be found in Appendix C.
@@
Effectiveness of DISC module.
We conducted experiments using the vanilla BERT without fine-tuning, initialized with bert-base-chinese, as shown in Table 6. Obviously, the general language model, without any fine-tuning or error correction strategies, cannot be applied to error correction tasks. However, after adding our DISC module, the performance of vanilla BERT improves significantly, giving the model basic error correction capabilities.
To further demonstrate the superiority of our method, we degrade the DISC module to a simple confusion set constraint decoding strategy. We investigate two confusion sets: one derived from our similarity computation strategy [^5] We treat a character pair as confused if their similarity score exceeds 0.5. and another pre-existing one provided by Wang et al. (2018). The results are shown in the second part of Table 6. From the results, we can see that both confusion sets fail to consistently improve performance, indicating the strategy’s sensitivity to confusion set quality. The confusion set from Wang et al. (2018) improves SIGHAN15 by covering over 99% of its erroneous pairs but degrades performance on other test sets, highlighting the domain-specific limitations of such confusion sets.
Effectiveness of components of the DISC module.
Analysis on Different Error Types.
Based on the similarity calculation strategy proposed in this paper, we separately filter the phonetic and glyph error types. Taking the phonetic error type as an example, the specific approach is as follows: for all edit pairs with a phonetic similarity < 0.5, we modify the source character to the golden character, resulting in a test set containing only phonetic edit pairs. The results are shown in Table 7.
According to our classification rule, the average proportions of phonetic and glyph on SIGHAN15 are 90.0% and 40.7%. From the results, it can be seen that DISC significantly improves performance on all error types. We find that, compared to phonetic errors, the original model exhibited a higher Recall metric and lower Precision metric on glyph errors, which makes the improvement brought by the DISC module more pronounced.
{NiceTabular}
[…]
$\mathrm{P}$ $\mathrm{R}$ $\mathrm{F}_{1}$
[…]
vanilla BERT 2.3 4.4 3.1 91.1
 + DISC 71.5 25.5 37.6 2.1
[…]
vanilla BERT 3.6 5.8 4.4 73.0
 + DISC 79.9 18.3 29.8 1.0
[…]
vanilla BERT 2.7 4.8 3.5 77.2
 + DISC 88.4 18.5 30.5 0.3
Table 5: Sentence-level performance of vanilla BERT.
Impact on Decoding Efficiency.
@@ 7 Related Work
These works fall into three categories. (1) At only the inference phase. Wang et al. (2019) and Bao et al. (2020) use the confusion set as constraints upon the search space, i.e., allowing the model to only consider characters in the confusion set.
(2) For data synthesis. Liu et al. (2021) use a confusion set $\mathcal{C}$ to synthesize data for training CSC models. For a given correct sentence, they randomly select a character (e.g., $c_{i}$), and replace it with an incorrect character (e.g., $c^{\prime}$). The replacement is constrained such that only pairs contained in the confusion set are considered, i.e., $(c_{i},c^{\prime})\in\mathcal{C}$.
@@
{NiceTabular}
[…]
$\mathrm{P}$ $\mathrm{R}$ $\mathrm{F}_{1}$
[…]
ReLM 65.2 77.6 70.9 19.3
 + DISC 69.8 78.0 73.7 14.3
[…]
ReLM 44.7 81.5 57.7 19.6
 + DISC 51.8 83.4 63.9 15.1
Table 7: Sentence-level performance of different error types on SIGHAN15.
@@
Besides the use of confusion sets, there exist some works that directly utilize phonetic and glyph information to enhance CSC models. Liu et al. (2021) and Li et al. (2022) add an extra task of predicting the phonetic of each input character. Xu et al. (2021) use GRU to encode Pinyin, and use CNN to encode glyphs (font pictures) for each input character, as extra character representations.
@@
Gou and Chen (2021) extract features such as probability and rank of the original character and the top 1 candidate character, and use SVM to judge modification retention. Yin et al. (2024) retrieve similar segments from the training set, and intervene in the decoding process based on the segment (n-gram) similarity between the retrieved segments and the input. Lv et al. (2023) employ a word dictionary in the target domain to assist the decoding process.
[…]
ReaLiSe 24.5
+ DISC 27.5 1.143$\times$
SCOPE 138.6
+ DISC 143.4 1.035$\times$
ReLM 12.7
+ DISC 12.8 1.010$\times$
@@ 8 Conclusions
Acknowledgements
We are deeply grateful to the anonymous reviewers for their thoughtful comments and dedicated efforts, which have greatly contributed to enhancing the quality and clarity of our work.
This work was supported by National Natural Science Foundation of China (Grant No. 62036004 and 62176173), Alibaba Group through Alibaba Innovative Research Program, and Project Funded by the Priority Academic Program Development of Jiangsu Higher Education Institutions.
@@ Appendix B Details of Datasets
Table 9: Statistics of the datasets.
@@ Appendix C Copy Punishment
Appendix D Prompt Example
@@
Figure 4: Prompt template used in GPT3.5 and GPT4.
In this work, we use the prompt-based method to activate the CSC ability of the GPT3.5 and GPT4. The prompt is shown in Figure 4.
Appendix E Experiments on Cleaned ECSpell
{NiceTabular}
[…]
$\mathrm{P}$ $\mathrm{R}$ $\mathrm{F}_{1}$
[…]
GPT4 62.0 62.0 62.0 7.3
ReLM 66.1 71.0 70.0 8.6
 + DISC 73.7 70.2 71.9 5.7
[…]
GPT4 45.1 57.5 50.6 24.8
ReLM 67.4 70.2 68.8 7.0
 + DISC 78.2 71.6 74.8 4.6
[…]
GPT4 71.7 67.6 69.5 1.7
ReLM 78.2 76.4 77.3 4.1
 + DISC 81.8 76.7 79.2 2.5
We discovered a serious data leakage issue in ECSpell. Specifically, the same correct sentence may appear multiple times in both the training and test sets, with only the location and type of errors varying. These sentences respectively account for 52.7%, 19.3%, and 28.2% of the ECSpell-LAW/MED/ODW training sets. We cleaned these duplicate sentences and reorganized the experiments, as shown in Table E.
Table 11: Sentence-level performance of LLMs, ReLM, and ReLM + DISC on the test sets of cleaned ECSpell.
The experimental results show that, compared to using the uncleaned training sets, ReLM’s performance on ECSpell significantly decreased, but it still outperforms GPT4. DISC also achieves stable performance improvement, with an impressive 6.0 F1 value increase on ECSpell-MED.
Appendix F Detailed Results
@@ D-P D-R D-F D-P D-R D-F D-P D-R D-F
In addition to the correction-level performance, we also present the detection-level experimental results of the CSC models, as shown in Table F. SCOPE + DR-CSC performs well at the detection level, primarily because they incorporate an additional detection network.
Since SIGHANs contain a lot of noise, we also conduct experiments on their revised versions (referred to as SIGHANs (rev.)) released by Yang et al. (2023b), which have undergone manual verification and error correction to ensure higher data quality. As shown in Table F and Table F, our DISC module also achieves consistent performance improvements on SIGHANs (rev.).